How much 12% acid must be added to 2 gallons of pure water to get a mixture that is 7%.
Answer by stanbon(75887)

How much 12% acid must be added to 2 gallons
of pure water to get a mixture that is 7%.
---
Equation::
acid + acid = acid
0*2 + 0.12x = 0.07(2+x)
------
0.12x = 0.14 + 0.07x
------
0.05x = 0.14
------
x = 14/5
x = 2.8 gallons (amt. of 12% solution to add)
----------------
Cheers,
Stan H.
------------